# Scheduling Your LLM Reinforcement Learning with Reasoning Trees

### v1

@@ 1 Introduction
Advancing the complex reasoning capabilities of Large Language Models (LLMs) remains a significant challenge, particularly in domains like mathematical problem-solving. Reinforcement Learning with Verifiable Reward (RLVR) (Gao et al., 2024; DeepSeek-AI et al., 2025), especially through policy optimization methods like GRPO (Shao et al., 2024), has emerged as a powerful paradigm to address this challenge. As shown in Figure 1 (a), in this framework, the space of potential solution paths for a query can be modeled as a specific ‘Reasoning Tree’ (Wang et al., 2025; Yang et al., 2025), where each node represents an intermediate reasoning step and each path represents a potential solution trajectory. From this perspective, RLVR operates as a dynamic ‘node-editing’ process of the reasoning tree: by rewarding correct paths and penalizing incorrect ones, the model iteratively refines its decision policy at each tree node. This optimization process gradually prunes branches that lead to low-quality or incorrect solutions, thereby improving overall reasoning accuracy.
In this paradigm, data scheduling plays a critical role in model performance (Hu et al., 2025; Yu et al., 2025; Li et al., 2025). The concept of data scheduling originates from curriculum learning (Bengio et al., 2009), which posits that models learn more effectively when training examples (queries) are organized in a meaningful sequence. Existing data scheduling strategies typically pre-define a‘difficulty’ metric for queries, and and schedule them from easy to hard to improve data efficiency and final performance (Xi et al., 2024; Chen et al., 2025b; a; Dai et al., 2025) However, from a reasoning tree perspective, current difficulty measure strategies exhibits a critical limitation: current methods estimate difficulty primarily via final solution accuracy, overlooking richer query-level characteristics such as the structural complexity of the reasoning tree. Accuracy alone is insufficient — low accuracy does not necessarily indicate that a query is inherently hard, and high accuracy does not guarantee ease of optimization. This inconsistency can undermine the efficacy of accuracy-based scheduling approaches. We illustrate this issue with the following examples.
@@ 2.1 Reinforcement learning with verifiable rewards in LLMs
Reinforcement learning with verifiable reward (RLVR), where the reward is computed by a rule-based verification function, has been shown to be effective in improving the reasoning capabilities of LLMs (Gao et al., 2024; DeepSeek-AI et al., 2025; Kimi et al., 2025; Zeng et al., 2025; Wen et al., 2025; Song et al., 2025). Typically, RLVR frameworks assign a binary reward by comparing the model’s generated output against a ground-truth solution, indicating whether it is correct or incorrect. This reward design obviates the need for complex outcome-based or process-based reward models, offering a straightforward yet potent approach. Recent advancements in policy optimization algorithms, such as PPO and GRPO, have further refined this paradigm (Schulman et al., 2017; Kazemnejad et al., 2024; Yuan et al., 2025; Yue et al., 2025; Shao et al., 2024; Yu et al., 2025; Liu et al., 2025; Zhang et al., 2025; Hu, 2025). In contrast to these studies, which focus on algorithmic improvements, our work builds upon the standard GRPO framework with a primary focus on designing a more effective training data schedule.
@@ 2.2 Data Scheduling Algorithm in LLM Reinforcement Learning
Various data scheduling strategies have been proposed to enhance the reasoning capabilities in LLM Reinforcement Learning. These can be broadly categorized into static selection and dynamic adjustment methods. Representative of static selection is LIMR (Li et al., 2025), which selected 1.4k examples from an 8.5k set for RLVR to match the performance of using the full set. In contrast, dynamic strategies make real-time adjustments during training. For instance, $R^{3}$ employs reverse curriculum reinforcement learning to simplify the model’s exploration space (Xi et al., 2024). LPPO dynamically adjusts data weights based on model accuracy and provides online ”prompts” for challenging problems to guide exploration (Chen et al., 2025b). SEED-GRPO focuses on model uncertainty (Chen et al., 2025a), using token entropy to dynamically re-weight data. Furthermore, DELT leverages training gradients to measure the quality and learnability of data (Dai et al., 2025), subsequently adjusting sample weights. However, these existing methods primarily rely on metrics such as accuracy and diversity, which overlook the structural relationships among data samples. In contrast, our work aims to quantify the learning potential of a query by analyzing its reasoning tree.
@@ 3.1 Group Relative Policy Optimization
The objective of the GRPO algorithm is to optimize a policy $\pi_{\theta}$ based on a group of generated responses (Shao et al., 2024; Yu et al., 2025). For a query $q$ from a dataset $\mathcal{D}$, the policy generates $G$ responses $\{o_{i}\}_{i=1}^{G}$. The token-level objective function is formulated as:
@@ 4 Motivation
The premise of this work is that path-based metrics such as accuracy are poor indicators of a query’s true learning difficulty. To illustrate our point, we supplement the example from the introduction with an experiment. As shown in Figure 2, we selected two distinct sets of 100 queries each from the DAPA-Math-17K dataset, using the Qwen2.5-Math-7B model. The blue line represents queries that have high initial accuracy but require extensive training to show further improvement. In contrast, the red line, standing for queries with simple tree structures, begins with low accuracy but improves rapidly. Although queries for the blue line appear ‘easier’ based on their initial accuracy, queries for the red line demonstrate greater potential. This discrepancy highlights that path-based metrics, like accuracy, are biased measurements for learning difficulty. This finding motivates us to design a new metric based on the structure of the reasoning tree.
@@ 5 Method
Figure 3: Overview of the Reasoning Tree Schedule (Re-Schedule) Algorithm.(a) Tree Construction: For each query, an approximate reasoning tree is constructed by sampling multiple solution paths from a base model. (b) R-Score Calculation: The tree’s structure is analyzed to compute the r-score, a metric quantifying the query’s learning potential. (c) Dynamic Weighting: The r-scores are used to dynamically weight each query during training, forming a curriculum that progresses from structurally simple (easy) to complex (hard) examples.
@@ 7.1 RL training Setups
Training setting We conduct experiments on two different models, including Qwen2.5-Math-7B and Qwen2.5-7B. We adapt our training codebase from verl (Sheng et al., 2025) and follow the training recipe of standard GRPO. Our training data is DAPO-Math-17k (Yu et al., 2025), containing only math problems with integer ground-truth answers. Both the KL-divergence and entropy loss terms are removed in our experiments. Generation batch size is set to $512$. Training is performed with top-p value of $1.0$ and temperature =1.0=1.0.
Evulation We evaluated our models and baselines on six widely used mathematical reasoning benchmarks: AIME24, AIME25, AMC23 (Li et al., 2024), MATH-500 (Hendrycks et al., 2021), Minerva Math (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). Validation is performed with a top-p value of $0.7$ and temperature =1.0=1.0 across all models and test sets. We use Math-Verify for training, validation, and final evaluation.
Baselines For the throughout comparison, we compare our method against 7 baselines, including standard GRPO (Shao et al., 2024), SimpleRL-Zoo (Zeng et al., 2025), Eurus-PRIME(Cui et al., 2025), OPO (Hao et al., 2025), ACC (curriculum learning based on accuracy, using sigmoid weighting), LPPO (Chen et al., 2025b), and Seed-GRPO (Chen et al., 2025a).
Our Methods Re-Schedule is implemented with two weighting schemes: ‘linear’ and ‘sigmoid’. Unless otherwise specified, the reasoning trees in our experiments are constructed with a branching factor of $k=4$, a maximum depth of $d=4$, and a token interval of $l=200$. The weighting schemes are defined as follows: 1. The ‘linear’ scheme uses $\gamma(t)=t/T$; 2. The ‘sigmoid’ scheme uses $\gamma(t)=\sigma\left(\left(\frac{t}{T}-0.5\right)\right)$. For both, we set the total number of epochs $T=10$.
Details of the training setup can be found in the Appendix B.
@@ 7.3 Ablation Experiment
We investigate the impact of the reasoning tree’s structure by varying the branching factor $k$ and maximum depth $d$. The choice of these parameters determines the fidelity of the approximated reasoning tree. While larger values for $k$ and $d$ theoretically provide a more accurate approximation and thus a more effective r-score, they also introduce a significant computational overhead. As shown in Table 3, our default configuration of $k=4$ and $d=4$ yields the best average performance (48.3%).
@@
We analyze the sensitivity of our method to the minimum $\omega_{\text{min}}$ and maximum $\omega_{\text{max}}$ weight hyperparameters, which control the dynamic range of the curriculum. Results in Table 4 show that our default setting of $\omega_{\text{min}}=0.5$ and $\omega_{\text{max}}=2.0$ achieves the highest average score (48.5). Decreasing the dynamic range by either reducing $\omega_{\text{max}}$ (to 1.5) or increasing $\omega_{\text{min}}$ (to 0.8) leads to performance degradation. This indicates that a sufficiently large weighting range is crucial for the curriculum to effectively differentiate between easy and hard samples. Conversely, an overly extreme range (e.g., $\omega_{\text{min}}=0.2$) also degrades performance, possibly because the curriculum excessively under-weights difficult queries. By assigning them a minimal weight for a prolonged period, the model is prevented from learning difficult queries. Furthermore, for additional experiments on the design choices for the r-score calculation, please see Appendix C.
@@ Appendix A Usage of LLMs
Appendix B Details of Experimental Setup
@@
Appendix C Supplementary Experiment: Effect of Metric Selection
@@
Re-Schedule (linear) 34.2 15.6 72.4 81.2 36.4 42.5 47.1
Re-Schedule (sigmoid) 34.2 16.0 71.1 81.8 42.3 44.4 48.3
@@
Re-Schedule (linear) 35.7 14.6 73.7 81.0 34.9 41.6 46.9
Re-Schedule (sigmoid) 33.1 16.7 71.1 82.0 39.0 42.4 47.4

### v2

@@ 1 Introduction
Advancing the complex reasoning capabilities of Large Language Models (LLMs) remains a significant challenge, particularly in domains like mathematical problem-solving. Reinforcement Learning with Verifiable Reward (RLVR) (1.3_RLVR; 1.1_deepseek_r1), especially through policy optimization methods like GRPO (1.6_grpo), has emerged as a powerful paradigm to address this challenge. As shown in Figure 1 (a), in this framework, the space of potential solution paths for a query can be modeled as a specific ‘Reasoning Tree’ (wang2025beyond; yang2025treerpo), where each node represents an intermediate reasoning step and each path represents a potential solution trajectory. From this perspective, RLVR operates as a dynamic ‘node-editing’ process of the reasoning tree: by rewarding correct paths and penalizing incorrect ones, the model iteratively refines its decision policy at each tree node. This optimization process gradually prunes branches that lead to low-quality or incorrect solutions, thereby improving overall reasoning accuracy.
In this paradigm, data scheduling plays a critical role in model performance (1.13_data_orz; 1.15_data_limr; yang2026whyattentionpatterns; renuncertainty). The concept of data scheduling originates from curriculum learning (bengio2009curriculum), which posits that models learn more effectively when training examples (queries) are organized in a meaningful sequence. Existing data scheduling strategies typically pre-define a‘difficulty’ metric for queries, and and schedule them from easy to hard to improve data efficiency and final performance (r3_add; chen2025data; chen2025seed; dai2025data) However, from a reasoning tree perspective, current difficulty measure strategies exhibits a critical limitation: current methods estimate difficulty primarily via final solution accuracy, overlooking richer query-level characteristics such as the structural complexity of the reasoning tree. Accuracy alone is insufficient — low accuracy does not necessarily indicate that a query is inherently hard, and high accuracy does not guarantee ease of optimization. This inconsistency can undermine the efficacy of accuracy-based scheduling approaches. We illustrate this issue with the following examples.
@@ 2.1 Reinforcement learning with verifiable rewards in LLMs
Reinforcement learning with verifiable reward (RLVR), where the reward is computed by a rule-based verification function, has been shown to be effective in improving the reasoning capabilities of LLMs (1.3_RLVR; 1.1_deepseek_r1; 1.2_kimi_k1.5; 2.2_rlvr; 2.3_rlvr_cur; 2.4_rlvr_cur_FastCuRL; wei-etal-2025-flexora; wei-etal-2025-paft; wei2025vflow; wei2025modelgen; maperturbation). Typically, RLVR frameworks assign a binary reward by comparing the model’s generated output against a ground-truth solution, indicating whether it is correct or incorrect. This reward design obviates the need for complex outcome-based or process-based reward models, offering a straightforward yet potent approach. Recent advancements in policy optimization algorithms, such as PPO and GRPO, have further refined this paradigm (1.5_ppo; 2.5_2_vineppo; 1.19_vc_ppo; 2.5_vapo_rlvr; 1.6_grpo; 1.10_dapo; 1.9_UnderstandingR1-Zero-LikeTraining_drgrpo; 1.7_SRPO; 1.12_REINFORCE++; hao2025rethinking; wei2025redit; liu2026automated; yang2025attentionpredictor). In contrast to these studies, which focus on algorithmic improvements, our work builds upon the standard GRPO framework with a primary focus on designing a more effective training data schedule.
@@ 2.2 Data Scheduling Algorithm in LLM Reinforcement Learning
Various data scheduling strategies have been proposed to enhance the reasoning capabilities in LLM Reinforcement Learning. These can be broadly categorized into static selection and dynamic adjustment methods. Representative of static selection is LIMR (1.15_data_limr), which selected 1.4k examples from an 8.5k set for RLVR to match the performance of using the full set. In contrast, dynamic strategies make real-time adjustments during training. For instance, $R^{3}$ employs reverse curriculum reinforcement learning to simplify the model’s exploration space (r3_add). LPPO (chen2025data) utilize the gradient of accuracy to prioritize data, effectively treating learning difficulty as a derivative of performance. Similarly, Seed-GRPO (chen2025seed) employs semantic diversity (uncertainty) as a proxy for difficulty. Furthermore, DELT leverages training gradients to measure the quality and learnability of data (dai2025data), subsequently adjusting sample weights. However, these methods rely on outcome-based proxies (e.g., accuracy), effectively treating reasoning as a flat sequence (zhao2025redone; zhang2025aligning). They overlook the inherent tree-structured solution space of reasoning tasks. In contrast, our approach explicitly leverages this topological structure. By analyzing the Reasoning Tree, we directly quantify a query’s ‘structural learnability’, providing a more precise and principled measure of difficulty than performance statistics alone.
@@ 3.1 Group Relative Policy Optimization
The objective of the GRPO algorithm is to optimize a policy $\pi_{\theta}$ based on a group of generated responses (1.6_grpo; 1.10_dapo). For a query $q$ from a dataset $\mathcal{D}$, the policy generates $G$ responses $\{o_{i}\}_{i=1}^{G}$. The token-level objective function is formulated as:
@@ 4 Motivation
The premise of this work is that path-based metrics such as accuracy are poor indicators of a query’s true learning difficulty. To illustrate our point, we supplement the example from the introduction with an experiment. As shown in Figure 2, we selected two distinct sets of 100 queries each from the DAPA-Math-17K dataset, using the Qwen2.5-Math-7B model. The blue line represents ‘Stagnant Samples’—queries with high initial accuracy but complex reasoning structures (low r-score). Their flat learning curve indicates that despite high initial performance, they are difficult to improve further. In contrast, the red line represents ‘Potential Samples’—queries with low initial accuracy but simple tree structures (high r-score). Their steep learning curve demonstrates high learnability, where a small amount of training yields significant gains. This discrepancy highlights that path-based metrics, like accuracy, are biased measurements for learning difficulty. This finding motivates us to design a new metric based on the structure of the reasoning tree.
@@ 5 Method
Figure 3: Overview of the Reasoning Tree Schedule (Re-Schedule) Algorithm.(a) Tree Construction: For each query, an approximate reasoning tree is constructed by sampling multiple solution paths from a base model (Note: This figureis for illustrative purposes only; our experiments use a tree with a depth of 4 and a width of 4, i.e., $k=4,d=4$.). (b) R-Score Calculation: The tree’s structure is analyzed to compute the r-score, a metric quantifying the query’s learning potential. (c) Dynamic Weighting: The r-scores are used to dynamically weight each query during training, forming a curriculum that progresses from structurally simple (easy) to complex (hard) examples.
@@ 7.1 RL training Setups
Training setting We conduct experiments on two different models, including Qwen2.5-Math-7B and Qwen2.5-7B. We adapt our training codebase from verl (sheng2025hybridflow) and follow the training recipe of standard GRPO. Our training data is DAPO-Math-17k (1.10_dapo), containing only math problems with integer ground-truth answers. Both the KL-divergence and entropy loss terms are removed in our experiments (hao2025rethinking). Generation batch size is set to $512$. Training is performed with top-p value of $1.0$ and temperature =1.0=1.0.
Evulation We evaluated our models and baselines on six widely used mathematical reasoning benchmarks: AIME24, AIME25, AMC23 (4_11_Numinamath), MATH-500 (4_3_math_dataset), Minerva Math (4_4_Minerva_dataset), and OlympiadBench (4_5_OlympiadBench). Validation is performed with a top-p value of $0.7$ and temperature =1.0=1.0 across all models and test sets. We use Math-Verify for training, validation, and final evaluation. We report avg@32 for all datasets. All results are presented as percentages.
Baselines For the throughout comparison, we compare our method against 7 baselines, including standard GRPO (1.6_grpo), SimpleRL-Zoo (2.2_rlvr), Eurus-PRIME(4_10_Eurus_prime), OPO (hao2025policy), ACC (curriculum learning based on accuracy, using sigmoid weighting), LPPO (chen2025data), and Seed-GRPO (chen2025seed).
Our Methods Re-Schedule is implemented with two weighting schemes: ‘linear’ and ‘sigmoid’. Unless otherwise specified, the reasoning trees in our experiments are constructed with a branching factor of $k=4$, a maximum depth of $d=4$, and a token interval of $l=200$. The weighting schemes are defined as follows: 1. The ‘linear’ scheme uses $\gamma(t)=t/T$; 2. The ‘sigmoid’ scheme uses $\gamma(t)=\sigma\left(\left(\frac{t}{T}-0.5\right)\right)$. For both, we set the total number of epochs $T=10$. Details of the training setup can be found in the Appendix C.
@@ 7.3 Ablation Experiment
We investigate the impact of the reasoning tree’s structure by varying the branching factor $k$ and maximum depth $d$. The choice of these parameters determines the fidelity of the approximated reasoning tree. While larger values for $k$ and $d$ theoretically provide a more accurate approximation and thus a more effective r-score, they also introduce a significant computational overhead. As shown in Table 3, our default configuration of $k=4$ and $d=4$ yields the best average performance (48.3%). For more detailed analysis, please see Appendices D.3 and D.4.
@@
We analyze the sensitivity of our method to the minimum $\omega_{\text{min}}$ and maximum $\omega_{\text{max}}$ weight hyperparameters, which control the dynamic range of the curriculum. Results in Table 4 show that our default setting of $\omega_{\text{min}}=0.5$ and $\omega_{\text{max}}=2.0$ achieves the highest average score (48.5). Decreasing the dynamic range by either reducing $\omega_{\text{max}}$ (to 1.5) or increasing $\omega_{\text{min}}$ (to 0.8) leads to performance degradation. This indicates that a sufficiently large weighting range is crucial for the curriculum to effectively differentiate between easy and hard samples. Conversely, an overly extreme range (e.g., $\omega_{\text{min}}=0.2$) also degrades performance, possibly because the curriculum excessively under-weights difficult queries. By assigning them a minimal weight for a prolonged period, the model is prevented from learning difficult queries. Furthermore, for additional experiments on the design choices for the r-score calculation, please see Appendix D.1.
Table 5: Computational cost vs. Performance gain. “Additional Cost” is relative to the total training time.
Tree Size ($k^{d}$) $\mathbf{3^{3}}$ $\mathbf{4^{3}}$ $\mathbf{3^{4}}$ $\mathbf{4^{4}}$ (Default)
Time Cost (hours) 3.48 6.21 6.70 22.67
Additional Cost +7.45% +13.30% +14.35% +48.54%
Avg Performance Gain +3.2 +3.0 +3.2 +4.0
7.4 Analysis Experiments
Figure 5: Average performance gain versus reasoning tree size and computational cost.
7.4.1 Computational Cost Analysis
As shown in Figure 5, we analyzed the trade-off between the offline tree construction cost and the resulting performance gain.
Table 5 presents the time cost measured on 8 $\times$ H20 GPUs. While larger trees ($4^{4}$) incur higher preprocessing costs compared to smaller trees ($3^{3}$), the cost remains manageable relative to the total training time (approx. 46 hours for 5 epochs), and the performance gains are substantial.
7.4.2 Impact of Ordering
Table 6: Comparison between Re-Schedule (Easy-to-Hard) and Reverse Schedule (Hard-to-Easy).
Schedule AIME24 AIME25 AMC23 MATH500 Minerva Olympiad Avg
Linear Mapping
Re-Schedule (Ours) 34.2 15.6 72.4 81.2 36.4 42.5 47.1
Reverse Schedule 31.9 14.0 67.6 81.0 37.8 41.8 45.7
Sigmoid Mapping
Re-Schedule (Ours) 34.2 16.0 71.1 81.8 42.3 44.4 48.3
Reverse Schedule 30.2 15.4 67.1 80.6 34.9 40.2 44.7
To validate the “easy-to-hard” curriculum design, we compared our method against a “Reverse Schedule” where lower r-score (harder) samples are prioritized first. As shown in Table 6, the Reverse Schedule leads to a significant drop in performance, confirming that starting with structurally simple samples is crucial for effective learning.
@@ Appendix A Usage of LLMs
Appendix B Acknowledgements
The authors would like to thank all the anonymous reviewers for their insightful comments and valuable suggestions. This research is supported by Anhui Provincial Natural Science Foundation (Grant No.2408085QF214), the Fundamental Research Funds for the Central Universities (Grant No.WK2080000206) and the Opening Project of the State Key Laboratory of General Artificial Intelligence (Project No.SKLAGI2025OP06). The work is also sponsored by the CodeBuddy (https://www.codebuddy.ai/). Special thanks go to Yi Liu and the team at Tencent for their generous support and assistance.
Appendix C Details of Experimental Setup
@@
Appendix D Supplementary Experiment
D.1 Effect of Metric Selection
@@
$\text{Re-Schedule}_{linear}$ 34.2 15.6 72.4 81.2 36.4 42.5 47.1
$\text{Re-Schedule}_{sigmoid}$ 34.2 16.0 71.1 81.8 42.3 44.4 48.3
@@
$\text{Re-Schedule}_{linear}$ 35.7 14.6 73.7 81.0 34.9 41.6 46.9
$\text{Re-Schedule}_{sigmoid}$ 33.1 16.7 71.1 82.0 39.0 42.4 47.4
@@
D.2 Variance Analysis
Table 8: Variance of Re-Schedule over multiple runs.
AIME24 AIME25 AMC23 MATH500 Minerva Olympiad
Avg. $\pm$ Var. 35.2 $\pm$ 0.1 16.0 $\pm$ 0.0 72.3 $\pm$ 0.4 82.2 $\pm$ 0.0 42.3 $\pm$ 0.6 44.4 $\pm$ 0.6
To assess the stability of our proposed Re-Schedule method, we conducted repeated experiments using different random seeds. Table 8 reports the variance. The results demonstrate that our method exhibits low variance. This confirms that the reported improvements are statistically stable and not due to random variation. Compared to performance improvements, the impact of variance is minimal.
D.3 Ablation on Token Interval $l$
To investigate the impact of $l$ on performance and its relationship with model capabilities, we conducted ablation studies on two base models: Qwen2.5-Math-7B and Qwen2.5-7B. We utilized the Sigmoid weighting mapping for these experiments.
Table 9 summarizes the average accuracy across six benchmarks (AIME24, AIME25, AMC23, MATH500, Minerva, Olympiad) for varying token intervals $l\in\{200,400,600,1200\}$.
Table 9: Impact of Token Interval $l$ on Average Accuracy (Sigmoid Mapping).
Interval ($l$) Qwen2.5-Math-7B (Avg) Qwen2.5-7B (Avg)
$l=200$ 48.3 44.5
$l=400$ 48.6 43.2
$l=600$ 48.9 43.1
$l=1200$ 46.0 41.1
The results indicate that Re-Schedule is generally robust to the choice of $l$ within the range of $[200,600]$. For the specialized math model (Qwen2.5-Math-7B), performance remains high and stable as $l$ increases to 600. For the general model (Qwen2.5-7B), while $l=200$ yields the best results, the performance drop at $l=600$ is relatively contained. A significant performance drop is observed for both models when $l=1200$. This suggests that when the interval is too large, the approximated reasoning tree becomes too coarse to capture the critical branching points necessary for effective r-score estimation.
D.4 Sensitivity to Branching Factor $k$, Depth $d$ and Modification Budget $M$
We investigated the impact of the reasoning tree size on performance by varying the branching factor $k$ and depth $d$ on the Qwen2.5-Math-7B model.
Table 10: Ablation study on branching factor $k$ (with fixed $d=4$).
Setting AIME24 AIME25 AMC23 MATH500 Minerva Olympiad Avg
Linear Mapping
$k=3$ 32.5 15.2 74.0 81.8 36.3 41.0 46.8
$k=4$ (Default) 34.2 15.6 72.4 81.2 36.4 42.5 47.1
$k=5$ 35.1 18.1 77.4 81.8 37.8 44.6 49.1
Sigmoid Mapping
$k=3$ 32.0 15.2 74.2 81.6 38.6 43.1 47.5
$k=4$ (Default) 34.2 16.0 71.1 81.8 42.3 44.4 48.3
$k=5$ 36.4 17.0 75.5 81.2 40.6 43.4 49.0
Table 11: Ablation study on tree depth $d$ (with fixed $k=4$).
Setting AIME24 AIME25 AMC23 MATH500 Minerva Olympiad Avg
Linear Mapping
$d=2$ 31.1 15.3 74.2 81.8 38.2 42.2 47.1
$d=3$ 31.4 14.6 72.7 82.0 39.7 43.3 47.3
$d=4$ (Default) 34.2 15.6 72.4 81.2 36.4 42.5 47.1
Sigmoid Mapping
$d=2$ 31.9 14.8 74.5 81.6 37.2 42.4 47.1
$d=3$ 33.2 16.4 73.0 80.0 41.9 41.6 47.7
$d=4$ (Default) 34.2 16.0 71.1 81.8 42.3 44.4 48.3
Table 12: Ablation study on node modification budget $M$ (with fixed $k=4,d=4,l=200$).
Setting AIME24 AIME25 AMC23 MATH500 Minerva Olympiad Avg
$M=5$ 33.6 15.4 72.2 79.0 40.4 43.1 47.3
$M=10$ (Default) 34.2 16.0 71.1 81.8 42.3 44.4 48.3
$M=15$ 34.7 16.0 71.8 82.0 41.6 42.4 48.1
Varying branching factor $k$: Fixing $d=4$ and $l=200$, we tested $k\in\{3,4,5\}$. As shown in Table 10, increasing $k$ generally improves performance, suggesting that a denser tree captures the structural difficulty more accurately.
Varying tree depth $d$: Fixing $k=4$ and $l=200$, we tested $d\in\{2,3,4\}$. Table 11 shows that deeper trees provide a better estimation of the reasoning structure, leading to improved downstream performance.
Varying node modification budget $M$: Finally, we assess the stability of our method with respect to the node modification budget $M$. Fixing $k=4,d=4$, and $l=200$, we evaluated performance across $M\in\{5,10,15\}$. As presented in Table 12, the results are relatively robust to changes in this parameter. While the default setting of $M=10$ yields the optimal average accuracy, varying the budget between 5 and 15 results in no significant performance degradation, indicating that the r-score remains a reliable metric across different budget constraints.
D.5 Generalization to Different Model Architectures
To demonstrate the broad applicability of our method beyond the Qwen2.5 family, we conducted additional experiments on the Qwen3-4B-Base model.
Table 13: Performance comparison on Qwen3-4B-Base.
Model AIME24 AIME25 AMC23 MATH500 Minerva Olympiad Avg
GRPO 24.2 21.8 52.4 86.0 39.4 43.4 44.5
ACC 24.8 23.3 59.3 88.8 41.6 42.0 46.6
Re-Schedule (Ours) 27.6 26.9 57.6 89.8 43.5 47.4 48.8
As shown in Table 13, Re-Schedule consistently outperforms both the standard GRPO baseline and the accuracy-based curriculum (ACC) across all benchmarks. This confirms that the effectiveness of the r-score is not limited to specific model architectures or sizes.
D.6 Dynamic R-Score Calculation
To determine if the r-score should be updated as the model evolves, we compared our standard static approach (computed once before training) with a dynamic approach where the r-score is re-computed and weights are updated three times during the training process.
Table 14: Comparison between Static and Dynamic R-Score updates on Qwen2.5-Math-7B.
Method AIME24 AIME25 AMC23 MATH500 Minerva Olympiad Avg
Linear Mapping
Static (Default) 34.2 15.6 72.4 81.2 36.4 42.5 47.1
Dynamic (3 updates) 34.6 14.9 75.3 80.0 39.7 41.6 47.8
Sigmoid Mapping
Static (Default) 34.2 16.0 71.1 81.8 42.3 44.4 48.3
Dynamic (3 updates) 35.3 15.2 74.2 82.4 42.7 43.6 48.9
As shown in Table 14, the dynamic approach yields performance comparable to the static baseline (e.g., 48.9% vs. 48.3% for Sigmoid mapping). Given the substantial computational cost of re-generating reasoning trees during training, we conclude that the static r-score serves as a sufficient and efficient prior for guiding the curriculum.
D.7 Generalization to Code Generation
Table 15: Performance comparison on Code Generation (LiveCodeBench v5).
Method pass@1 pass@4
GRPO 25.4 35.4
$\text{ACC}_{sigmoid}$ 25.8 36.0
$\text{Re-Schedule}_{sigmoid}$ 26.3 37.8
To validate the generalization capability of Re-Schedule beyond mathematical reasoning, we extended our evaluation to the domain of Code Generation. We utilized DeepSeek-R1-Distill-Qwen-1.5B as the base model. The model was trained on the ArcherCodeR dataset wang2025stabilizing, which contains 6,753 code generation tasks. For evaluation, we used the LiveCodeBench v5 benchmark jain2024livecodebench. We report pass@1 and pass@4 metrics (averaged over 8 samples).
As shown in Table 15, Re-Schedule consistently outperforms both the standard GRPO baseline and the accuracy-based curriculum (ACC). Specifically, our method achieves a +0.9% improvement in pass@1 and a significant +2.4% improvement in pass@4 compared to GRPO. These results confirm that the structural insights captured by the r-score are effective in the coding domain, where the “reasoning tree” corresponds to the decision space of code logic and syntax. Looking forward, the tree-based structural metrics could also be adapted to improve multi-modal data generation pipelines and self-instructed compositional code captioning (wang2024world; wang2025grasp; wang2025vgr; hao2026recreate; wang2025traceable; lei2025scalability).